# A Closer Look at AUROC and AUPRC under Class Imbalance

### v2

@@ 2.1 Relationship Between AUROC and AUPRC
Let $f:\mathcal{X}\to(0,1)$ be a binary classification model over a data distribution $\mathsf{x},\mathsf{y}$ of model inputs $x\sim\mathsf{x},x\in\mathcal{X}$ and labels $y\sim\mathsf{y},y\in(0,1)$. Then,
@@ 2.2 AUPRC prioritizes high-score mistakes, AUROC treats all mistakes equally
Understanding how a given evaluation metric prioritizes the correction of various kinds of model mistakes or errors offers significant insight into when that metric should be used for optimization or model selection. To examine this topic for AUROC and AUPRC, consider the following definition of a model “mistake”:
@@
Let $f:\mathcal{X}\to(0,1)$ be a binary classification model over a domain $\mathcal{X}$ being evaluated over a static, finite dataset $\boldsymbol{X},\boldsymbol{y}=\{(x_{1},y_{1}),\ldots,(x_{N},y_{N})\}$, for $x_{i}\in\mathcal{X},y_{i}\in\{0,1\},$ and $N\in\mathbb{N}$. We assume for convenience that $f$ is an injective map and all $x_{i}$ are distinct (i.e., $\forall(i,j):f(x_{i})\neq f(x_{j})$).
We say that a model makes a mistake at samples $(x_{i},x_{j})$ if:
@@
Define $f,\mathcal{X},\boldsymbol{X},\boldsymbol{y}$ and $N$ as in Definition 2.1. Further, suppose without loss of generality that the dataset $\boldsymbol{X}$ is ordered such that $f(x_{i})<f(x_{i+1})$ for all $i$. Then, let us define $M=\{i|(x_{i},x_{i+1})\text{ is a \emph{mistake} for model $f$}\}$. Define $f_{i}^{\prime}$ to be a model that is identical to $f$ except that the probabilities assigned to $x_{i}$ and $x_{i+1}$ are swapped:
@@
The proof for Theorem 2 can be found in Appendix D. This proof simply stems from the fact that correcting a single mistake $(x_{i},x_{j})$ (as defined in Definition 2.1) always changes the false positive rate by the same amount, and only changes it at the threshold $f(x_{i})$. This, combined with the formalization of AUROC and AUPRC in Theorem 1, establishes the proof.
@@ 2.3 AUPRC is explicitly discriminatory in favor of high-scoring subpopulations
Let $f,\mathcal{X},\boldsymbol{X},\boldsymbol{y},N,M,$ and $f^{\prime}_{j}$ all be defined as in Theorem 2. Further, suppose that in this setting the domain $\mathcal{X}$ now contains an attribute defining two subgroups, $\mathcal{A}=\{0,1\}$, such that for any sample $(x_{i},y_{i})$, $a_{i}$ denotes the subgroup to which that sample belongs. Let $f$ be perfectly calibrated for samples in subgroup $a=0$, such that $p_{\mathsf{y}|\mathsf{a}=0,f(\mathsf{x})=t}(1)=t$. Let $p_{0}=p_{\mathsf{y}|\mathsf{a}=0}(1)$ denote the prevalence of the label $y$ over subgroup $a=0$. Then,
$\lim_{p_{0}\to 0}P\left(a_{i}=a_{i+1}=1\middle|i=\operatorname*{arg\,max}_{j\in\mathcal{M}}\left(\mathrm{AUPRC}(f_{j}^{\prime})\right)\right)=1.$
@@ 6 Limitations and Future Works
While our analyses are thorough and compelling, there are still a number of areas for further improvement and future work. Firstly, our theoretical findings can be refined and generalized to less restrictive settings, that take into account the difficulty of the target task (which may differ between subgroups) or does not require models to be calibrated (in the case of Theorem 3). Further, extending our real-world experiments to more fairness datasets and identifying more nuanced ways to probe the impact of metric choice on disparity measures would significantly strengthen this work. Lastly, These analyses can be extended to consider other metrics, such as the area under the precision-recall-gain curve (Flach & Kull, 2015), the area under the net benefit curve (Talluri & Shete, 2016; Pfohl et al., 2022), and single-threshold, deployment centric metrics as well.
@@ Appendix B Notation
Let $f_{\boldsymbol{\theta}}:\mathcal{X}\to\mathcal{(}0,1)$ denote a model, with $\mathcal{X}$ its input domain and $(0,1)=\{r\in\mathbb{R}|0<r<1\}$ its output domain for a binary classification task. Let the random variable $\mathsf{x}$ describe the distribution of input samples $x_{i}\in\mathcal{X}$, and define random variable $\mathsf{p}=f_{\boldsymbol{\theta}}(\mathsf{x})$ to be the distribution of scores output by the model over input samples and let $\mathsf{y}$ be the random variable describing labels $y_{i}\in\{0,1\}$ for the underlying learning task. Throughout the paper, $\boldsymbol{\theta}$ may be omitted if it is clear from context. We will occasionally also use the notation $\mathsf{p}_{+}$ and $\mathsf{p}_{-}$ to reflect the conditional distributions of model outputs conditioned on the label being 1 or 0, respectively:
@@ Appendix D Proof of Theorem 2
As AUROC weights the false positive rate at all positive samples equally and the false positive rate is proportional to the number of false positives, this shows that AUROC will improve by a constant amount no matter which atomic mistake is fixed. In contrast, as AUPRC weights false positives inversely by the model’s firing rate, it will improve by an amount that is directly linearly correlated with the inverse of the model’s firing rate, implying that it favors mistakes with higher scores and disfavors mistakes with lower scores. ∎
@@ F.3 Details on optimization procedures
    Sequentially Fixing Atomic Mistakes. We sequentially correct atomic mistakes, as defined in Figure 1. At each step, we first discover the set of all atomic mistakes $M$. To maximize AUROC, we randomly select a pair $(S_{i},S_{j})\in M$, and swap their scores in $S$, i.e. $S^{\prime}_{i}=S_{j},S^{\prime}_{j}=S_{i}$. To maximize AUPRC, we swap the scores for the pair $(S_{i},S_{j})=\arg\max_{(s_{i},s_{j})\in M}s_{j}$. We repeat this process for 50 steps, with each one sequentially fixing another atomic mistake in $S$. Results for this setting are shown in Figures 3(b) and 3(a).
@@
    Sequentially Permuting Nearby Scores. We first sort $S$ and $Y$ such that $S$ is in ascending order. We apply a random permutation to $S$ by re-indexing it using a random ordering, but such that scores are not shuffled too far from their original index. Let $\sigma$ be the ordered sequence $(1,2,...,n)$. Define $\Omega$ to be the set of all permutations of $\sigma$, such that for all $\omega\in\Omega$, $|\omega_{i}-\sigma_{i}|\leq\gamma$ for $i\in\{1,...,n\}$. At each step, we sample $\omega\in\Omega$ with $\gamma=3$ twenty times, where each $\omega$ corresponds to a new candidate ordering of $S$. We compute the selection metric for each of the twenty orderings, and return $S^{\prime}$ to be the score permutation that achieves the maximum value for the selection metric. We repeat this procedure for 25 steps, setting $S$ at each step to be the $S^{\prime}$ output from the previous step. Results for this setting are shown in Figures 3(d) and 3(c).
@@ H.4.1 Final Papers
Claim

### v3

@@ 2.1 Relationship Between AUROC and AUPRC
Let $\mathcal{X},\mathcal{Y}={0,1}$ represent a paired feature and binary classification label space from which i.i.d. samples $(x,y)\in\mathcal{X}\times\mathcal{Y}$ are drawn via the joint distribution over the random variables $\mathsf{x},\mathsf{y}$. Let $f:\mathcal{X}\to(0,1)$ be a binary classification model outputting continuous probability scores over this space. Then,
@@ 2.2 AUPRC prioritizes high-score mistakes, AUROC treats all mistakes equally
Understanding how a given evaluation metric prioritizes the correction of various kinds of model mistakes or errors offers significant insight into when that metric should be used for optimization or model selection. To examine this topic for AUROC and AUPRC, consider the following definition of an “incorrectly ranked adjacent pair”, which we will colloquially refer to as a “model mistake”:
@@
Let $f,\mathcal{X},\mathcal{Y},\mathsf{x},\mathsf{y}$ be defined as in Theorem 1. Further, let us suppose we have sampled a static dataset from $\mathsf{x},\mathsf{y}$ for evaluation which will be denoted $\boldsymbol{X},\boldsymbol{y}=\{(x_{1},y_{1}),\ldots,(x_{N},y_{N})\}$, for $x_{i}\in\mathcal{X},y_{i}\in\{0,1\},$ and $N\in\mathbb{N}$. We assume for convenience that $f$ is an injective map and all $x_{i}$ are distinct (i.e., $\forall(i,j):x_{i}\neq x_{j}$ which, by injectivity of $f$, implies that $f(x_{i})\neq f(x_{j})$).
We say that $(x_{i},x_{j})$ are an incorrectly ranked adjacent pair and thus that the model makes a “mistake” at samples $(x_{i},x_{j})$ if:
@@
Define $f,\mathcal{X},\boldsymbol{X},\boldsymbol{y}$ and $N$ as in Definition 2.1. Further, suppose without loss of generality that the dataset $\boldsymbol{X}$ is ordered such that $f(x_{i})<f(x_{i+1})$ for all $i$. Then, let us define $M=\{i|(x_{i},x_{i+1})\text{ is an \emph{incorrectly ranked adjacent pair} for model $f$}\}$. Define $f_{i}^{\prime}$ to be a model that is identical to $f$ except that the probabilities assigned to $x_{i}$ and $x_{i+1}$ are swapped:
@@
The proof for Theorem 2 can be found in Appendix D. This proof simply stems from the fact that correcting a single mistake $(x_{i},x_{j})$ (as defined in Definition 2.1) always changes the false positive rate by the same amount, and only changes it at the threshold $f(x_{i})$. This, combined with the formalization of AUROC and AUPRC in Theorem 1, establishes the proof. Note that this Theorem can be trivially extended to include a case where ties are possible simply by noting that “swapping” two samples $x_{i}$ and $x_{j}$ in the manner of the theorem results in no change to either AUROC or AUPRC, and similarly by the same reasoning separating any tie in the appropriate direction will improve AUROC uniformly over samples and will improve AUPRC in a manner monotonic with model score.
@@ 2.3 AUPRC is explicitly discriminatory in favor of high-scoring subpopulations
Let $f,\mathcal{X},\boldsymbol{X},\boldsymbol{y},N,M,$ and $f^{\prime}_{j}$ all be defined as in Theorem 2. Further, suppose that in this setting the domain $\mathcal{X}$ now contains an attribute defining two subgroups, $\mathcal{A}=\{0,1\}$, such that for any sample $(x_{i},y_{i})$, $a_{i}$ denotes the subgroup to which that sample belongs. Let $f$ be perfectly calibrated for samples in subgroup $a=0$, such that $P(y=1|a=0,f(x)=t)=t$. Let $p_{0}=p(y=1|a=0)$ denote the prevalence of the label $y$ over subgroup $a=0$. Then,
$\lim_{p_{0}\to 0}P\left(a_{i}=a_{i+1}=1\middle|i=\operatorname*{arg\,max}_{j\in M}\left(\mathrm{AUPRC}(f_{j}^{\prime})\right)\right)=1.$
@@ 6 Limitations and Future Works
While our analyses are thorough and compelling, there are still a number of areas for further improvement and future work. Firstly, our theoretical findings can be refined and generalized to less restrictive settings, that take into account the difficulty of the target task (which may differ between subgroups) or does not require models to be calibrated (in the case of Theorem 3). Further, extending our real-world experiments to more fairness datasets and identifying more nuanced ways to probe the impact of metric choice on disparity measures would significantly strengthen this work. Lastly, These analyses can be extended to consider other metrics, such as the area under the precision-recall-gain curve (Flach & Kull, 2015), the area under the net benefit curve (Talluri & Shete, 2016; Pfohl et al., 2022), and single-threshold, deployment centric metrics as well. In addition, further expanding theoretical analyses, such as examining how Theorem 3 can be softened to have less restrictive requirements regarding calibration or whether or not these results imply anything about the viability or safety of post-hoc calibration of models optimized either through AUPRC or AUROC would also be extremely interesting.
@@ 7 Conclusion
Acknowledgements
MBAM is supported by a Berkowitz Postdoctoral Fellowship. JG is funded by the National Institute of Health through DS-I Africa U54 TW012043-01 and Bridge2AI OT2OD032701.
@@ Appendix B Notation
Let $\mathcal{X},\mathcal{Y}={0,1}$ represent a paired feature and binary classification label space from which i.i.d. samples $(x,y)\in\mathcal{X}\times\mathcal{Y}$ are drawn via the joint distribution over the random variables $\mathsf{x},\mathsf{y}$. Let $f_{\boldsymbol{\theta}}:\mathcal{X}\to(0,1)$ be a binary classification model parametrized by $\boldsymbol{\theta}\in\mathbb{R}^{d}$ for some $d\in\mathbb{N}$ outputting continuous probability scores over this space.
We define random variable $\mathsf{p}=f_{\boldsymbol{\theta}}(\mathsf{x})$ to be the distribution of scores output by the model over input samples. Throughout the paper, $\boldsymbol{\theta}$ may be omitted if it is clear from context. We will occasionally also use the notation $\mathsf{p}_{+}$ and $\mathsf{p}_{-}$ to reflect the conditional distributions of model outputs conditioned on the label being 1 or 0, respectively:
@@ Appendix C Proof of Theorem 1
Recall that all notation is defined formally in Appendix B.
@@ Appendix D Proof of Theorem 2
As AUROC weights the false positive rate at all positive samples equally and the false positive rate is proportional to the number of false positives, this shows that AUROC will improve by a constant amount no matter which atomic mistake is fixed. In contrast, as AUPRC weights false positives inversely by the model’s firing rate, it will improve by an amount that is directly linearly correlated with the inverse of the model’s firing rate, implying that it favors mistakes with higher scores and disfavors mistakes with lower scores.
Note that as we use strict inequalities in our definition of the decision rule underlying the FPR here, a pair of scores that are tied but have different labels will not induce a false positive at the corresponding positively labeled sample’s threshold, so separating such ties will have no impact on AUROC whatsoever. It would similarly not impact AUPRC as neither the FPR nor the model firing rate will decrease when the negative sample within the tie is perturbed to be strictly below the positive sample. ∎
@@ F.3 Details on optimization procedures
    Sequentially Fixing Atomic Mistakes. We sequentially correct atomic mistakes, as defined in Figure 1. At each step, we first discover the set of all atomic mistakes $M$. To maximize AUROC, we randomly select a pair $(S_{i},S_{j})\in M$, and swap their scores in $S$, i.e. $S^{\prime}_{i}=S_{j},S^{\prime}_{j}=S_{i}$. To maximize AUPRC, we swap the scores for the pair $(S_{i},S_{j})=\arg\max_{(s_{i},s_{j})\in M}s_{j}$. We repeat this process for 50 steps, with each one sequentially fixing another atomic mistake in $S$. Results for this setting are shown in Figures 4(b) and 4(a).
@@
    Sequentially Permuting Nearby Scores. We first sort $S$ and $Y$ such that $S$ is in ascending order. We apply a random permutation to $S$ by re-indexing it using a random ordering, but such that scores are not shuffled too far from their original index. Let $\sigma$ be the ordered sequence $(1,2,...,n)$. Define $\Omega$ to be the set of all permutations of $\sigma$, such that for all $\omega\in\Omega$, $|\omega_{i}-\sigma_{i}|\leq\gamma$ for $i\in\{1,...,n\}$. At each step, we sample $\omega\in\Omega$ with $\gamma=3$ twenty times, where each $\omega$ corresponds to a new candidate ordering of $S$. We compute the selection metric for each of the twenty orderings, and return $S^{\prime}$ to be the score permutation that achieves the maximum value for the selection metric. We repeat this procedure for 25 steps, setting $S$ at each step to be the $S^{\prime}$ output from the previous step. Results for this setting are shown in Figures 4(d) and 4(c).
@@ H.4.1 Final Papers
Claim References Commentary
Precision-recall curves or other associated metrics may more appropriately reflect deployment objectives than the receiver operating characteristic. (Cook & Ramadas, 2020; Leisman, 2018; Yang et al., 2015; Muthukrishna et al., 2019; Deng et al., 2023; Harer et al., 2018; Ahmed & Courville, 2020) While this claim is true, the informativeness of the PR curve for target deployment metrics is not sufficient to conclude that the AUPRC is superior to the AUROC in all cases of class imbalance. Despite this, it is often taken to assert this more general claim without caveat.
AUPRC does not depend on the number of true-negatives, so will be less optimistic than the AUROC (Leisman, 2018; Kyono et al., 2018; Adler, 2021; Meister et al., 2022; Mosteiro et al., 2021; Showalter & Wu, 2019; Cranmer & Desmarais, 2016; Domingues et al., 2020; Rezvani et al., 2021; Hsu et al., 2020; Ju et al., 2018; Pashchenko et al., 2018; Romero et al., 2022; Vens et al., 2008) As shown in Theorem 1, AUROC and AUPRC can both be naturally expressed as a function of the expectation of the model’s false positive rate. More generally, lack of dependence on one quadrant among the mutually dependent four quadrants of a confusion matrix is not an informative property for the AUROC and AUPRC metrics.
AUPRC will often be significantly lower, farther from optimality, and/or will grow more non-linearly as model performance improves than AUROC for low-prevalence tasks (Leisman, 2018; Yang et al., 2015; Mehboudi et al., 2022; Cranmer & Desmarais, 2016) Metric utility for model comparison depends on how appropriately it prioritizes model improvements, and is therefore less about the raw magnitude of the metric and more about the situations in which the order of a set of models will differ under one metric vs. another. One could easily make AUROC yield smaller values or grow more quickly near optimality by simply exponentiating it, but this would not yield a better metric.
AUPRC depends on prevalence, which is a desirable property (Navarro et al., 2022) This statement is too vague to be formally evaluated; whether or not this dependence on prevalence is desirable depends on the context. For model comparison in general, we argue it is not desirable in this form as it induces the biases inhere in AUPRC previously discussed.
AUPRC better captures differentiating a positive sample with high score from a “hard” negative sample (“hard” meaning one also with high score) (Kiran et al., 2018) While this claim is true by Theorem 2, it is not clear why this would be desired in general; this implicitly favors comparing “hard” negatives against “easy” positives as opposed to “easy” negatives against “hard” positives.
AUROC is otherwise “optimistic” in low-prevalence settings (Cook & Ramadas, 2020; Afanasiev et al., 2021; Wu et al., 2020; Miao & Zhu, 2022; Cho et al., 2021; Hagedoorn & Spanakis, 2017; Yang et al., 2022a; Mangolin et al., 2022; Silva et al., 2022; Lyu et al., 2021; Ahmed & Courville, 2020; Zhang et al., 2021; Lund et al., 2019; Ruff et al., 2021; Chakraborty et al., 2023; Rajabi & He, 2021; Ando & Huang, 2017; Kulkarni et al., 2021; Karadzhov et al., 2022; Hibshman & Weninger, 2023) This claim is underspecified, and un-true. AUROC always means the same thing, probabilistically, and that meaning independent from class imbalance.
AUPRC focues more on the positive (minority) class (Tusfiqur et al., 2022; Wu et al., 2020; Si & Roberts, 2021; Narayanan et al., 2022; Babaei et al., 2021; Garcin & Stéphan, 2021; Deng et al., 2023; Alvarez et al., 2022; Mosteiro et al., 2021; Showalter & Wu, 2019; Pang et al., 2023; Ozyegen et al., 2022; Danesh Pazho et al., 2023; Kiran et al., 2018; López et al., 2013; Rao et al., 2022; Ntroumpogiannis et al., 2023) This is unfounded; both AUROC and AUPRC are weighted expectations over the model’s false positive rate—AUPRC cares more about samples in regions of low firing rate, not explicitly about positive or minority samples.
AUROC can not appropriately detect models with poor recall (Navarro et al., 2022) This claim is unfounded; the AUROC clearly depends on the model’s recall. Besides, if recall is the measure of interest, then that should be measured explicitly.
Table 2: Various arguments and our responses to them present for this claim in the literature.
Claim References Valid? Commentary
Precision-recall curves or other associated metrics may more appropriately reflect deployment objectives than the receiver operating characteristic. (Cook & Ramadas, 2020; Leisman, 2018; Saito & Rehmsmeier, 2015; Yuan et al., 2015; Bleakley et al., 2007; Ozenne et al., 2015; Rosenberg, 2022; Zhou et al., 2020; Lichtnwalter & Chawla, 2012; Yang et al., 2015) While this claim is true, the informativeness of the PR curve for target deployment metrics is insufficient to conclude that the AUPRC is superior to the AUROC in all cases of class imbalance. Despite this, it is often taken to assert this more general claim without caveat.
AUPRC does not depend on the number of true negatives, so will be less optimistic than the AUROC (Leisman, 2018; Goadrich et al., 2006; Cranmer & Desmarais, 2016) As shown in Theorem 1, AUROC and AUPRC can both be naturally expressed as a function of the expectation of the model’s false positive rate. More generally, the lack of dependence on one quadrant among the mutually dependent four quadrants of a confusion matrix is not an informative property for the AUROC and AUPRC metrics.
AUPRC will often be significantly lower, farther from optimality, and/or will grow more non-linearly as model performance improves than AUROC for low-prevalence tasks (Leisman, 2018; Yuan et al., 2015; Goadrich et al., 2006; Mazzanti, 2023; Rosenberg, 2022; Zhou et al., 2020; Lichtnwalter & Chawla, 2012; Yang et al., 2015; Cranmer & Desmarais, 2016) Metric utility for model comparison depends on how appropriately it prioritizes model improvements. Therefore, it is less about the raw magnitude of the metric and more about the situations in which the order of a set of models will differ under one metric vs. another. One could easily make AUROC yield smaller values or grow more quickly near optimality by simply exponentiating it, but this would not yield a better metric.
AUPRC depends on prevalence, which is a desirable property (Saito & Rehmsmeier, 2015; Goadrich et al., 2006; Yuan et al., 2015) This statement is too vague to be formally evaluated; whether or not this dependence on prevalence is desirable depends on the context. For model comparison in general, we argue it is not desirable in this form as it induces the biases in AUPRC previously discussed.
AUPRC better captures differentiating a positive sample with high score from a “hard” negative sample (“hard” meaning one also with high score) (Rosenberg, 2022) While this claim is true by Theorem 2, it is not clear why this would be desired in general; this implicitly favors comparing “hard” negatives against “easy” positives as opposed to “easy” negatives against “hard” positives.
Table 3: Various arguments and our responses to them present for this claim in the literature.